OpenStudy (aaronandyson):

The speed of a boat in still water is 15 km/hr.It can go 30 km upstream and return downstream to the original point in 4 hours 30 minutes.Find the speed of the stream.

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (welshfella):

if the speed of the stream is v km / hour speed of boat downstream = 15 + v and upstream speed = 15 - v distance = speed * time downstream 30 = t1(15 + v) upstream 30 = t2 (15 - v) where t1 and t2 are the times for downstream and upstream also we have t1 + t2 = 4.5

OpenStudy (welshfella):

so you need to solve the following system of equations t1 + t2 = 4.5 t1(15 + v) = t2(15 - v)

OpenStudy (welshfella):

I'm missing something here as we've got 3 unknown and only 2 equations

OpenStudy (welshfella):

sorry i have no time at the moment . Gotta go

OpenStudy (irishboy123):

\[\frac{30}{15 - v} + \frac{30}{15 + v} = \frac{9}{2} \ [hrs]\] \[\frac{30(15 + v) + 30(15 - v)}{15^2 - v^2} = \frac{9}{2}\] \[\frac{2(30)(15)}{15^2 - v^2} = \frac{9}{2}\] \[2(2)(30)(15) = 9(15^2 - v^2)\] etc,